particle moving with uniform acceleration along x axis at x1 = 10 and x2 = 20 at t = 6 and t = 10 respectively. If its velocity at t = 6 is 0.5 m/s it acceleration is
1
Expert's answer
2018-12-03T08:49:49-0500

Answer on Question # 83397, Mechanics/ Relativity

particle moving with uniform acceleration along x axis at x1 = 10 and x2 = 20 at t = 6 and t = 10 respectively. If its velocity at t = 6 is 0.5 m/s it acceleration is

Solution

v2 = v1 + at

v22 = v12 + 2aΔx

Then

(v1 +at)2 = v12 + 2aΔx (1)

where

Δx= x2 - x1 = 20-10= 10 m

Δt = t2 - t1 =10-6 = 4 s

v1 = 0.5 m/s

v2 = 0.5 + a*4

Solve equation (1)for a:

(0.5 + a*4)2 = (0.5)2 + 2a*10

(0.5 + a*4)2 -(0.5)2 = 2a*10

(0.5+4a-0.5)(0.5+4a+0.5) =20a

4a(1+4a)=20a

1+4a =5

a=1 (m/s2)

Answer: a=1 (m/s2)